# Self-consistent neutron stars in a class of massive vector-tensor gravity

### v1

@@ I Introduction
and by imposing Eq. (1) throughout the entire vacuum region [19]. Subsequent studies have further shown that Einstein-bumblebee gravity admits exact black hole solutions involving a richer set of physical parameters, including electric charge [20], magnetic charge [21], and NUT charge [22]. This makes the theory a convenient and concrete framework for constraining spontaneous Lorentz symmetry breaking through astrophysical observations.
@@
Technically, there is no intrinsic difficulty in deriving compact star solutions from the Einstein-bumblebee gravity Lagrangian. Previous attempts failed primarily because of an unnecessarily strong and somewhat ad hoc theoretical assumption that condition (1) be imposed throughout the entire vacuum region. Einstein-bumblebee gravity, as a massive non-minimally coupled vector-tensor theory, is a higher-derivative gravity theory, for which vacuum spherically symmetric solutions are not unique. In this context, black hole solutions arise only as special vacuum solutions for which condition (1) can be consistently satisfied. For compact stars, however, different equations of state (EOSs) and central densities generically lead to different vacuum boundary conditions at the stellar surface. As a result, the exterior vacuum spacetime of a compact star may deviate from the equal-mass black hole geometry, and condition (1) is therefore no longer satisfied. Nevertheless, such deviations are confined mainly to the strong-field region, since both black hole and compact star solutions necessarily share the same boundary conditions at spatial infinity. This implies that assumption (1) need not be imposed a priori: it is naturally violated in the strong-field region while being dynamically enforced in the weak-field regime through asymptotic boundary conditions at spatial infinity. This treatment is well established in massive non-minimally coupled scalar-tensor gravity [29, 30, 31, 32]. Consequently, consistency with previously obtained black hole solutions and existing observational constraints is preserved, and earlier Solar System tests of spontaneous Lorentz symmetry breaking based on exact black hole spacetimes remain physically meaningful.
@@ III.2 Vacuum equations
where the subscripts “in” and “ext” denote the interior and exterior solutions, respectively. We then continue the integration outward through the vacuum region up to spatial infinity. The vacuum field equations are given by Eqs. (12)–(13) with $p=0$ and $\rho=0$. The zeroth-order vacuum equations can be organized into a coupled system nonlinear ODEs, which take the schematic form
@@ IV Equilibrium configurations and moments of inertia
To verify the consistency of the neutron star solutions obtained in this work, we first plot the metric functions ($f,\phi^{-1/2},\rho,\lambda^{\prime},w/\Omega$), as functions of the radial coordinate $r$ in Fig. 1, for representative values of the parameters $\ell$ and $\alpha$.
Figure 1: The plots illustrate the numerical solutions of the functions ($f,\phi^{-1/2},\rho,\lambda^{\prime},w/\Omega$) in massive vector-tensor gravity with the parameters $\alpha=10^{-4}\alpha_{\star}$ (solid lines) and $\alpha=10^{-2}\alpha_{\star}$ (dashed lines). The parameter $\ell$ is fixed to $\ell=10^{-10}$. The central density is set to $\rho_{0}=1.2\times 10^{15}\textup{g/cm}^{3}$, and the SLy EOS is adopted.
Motivated by the bounds $(10^{-12},10^{-9})$ on the parameter $\ell$ inferred from three classical Solar System tests [19], we choose $\ell=10^{-10}$ for our numerical analysis. We consider two values of the parameter $\alpha$ namely $\alpha=10^{-4}\alpha_{\star}$ and $\alpha=10^{-2}\alpha_{\star}$. The central density is fixed at $\rho_{0}=1.2\times 10^{15}\textup{g/cm}^{3}$. For both choices of $\alpha$, the resulting neutron star mass is $1.51M_{\odot}$, while the corresponding radii are $7.48r_{\star}$ and $7.58r_{\star}$, and the moments of inertia are $34.92I_{\star}$ and $34.25I_{\star}$. From the upper panel of Fig. 1, one can clearly see that $\phi^{2}f$ deviates from unity both inside the neutron star and in the exterior strong-field region, while it approaches unity only in the weak-field regime and asymptotically at spatial infinity, as required by the boundary condition (24). Moreover, the deviation becomes more pronounced as the potential parameter $\alpha$ decreases.
Next, we plot the mass-radius $M$-$R$ and mass-central density $M$-$\rho_{0}$ relations for the SLy EOS over a range of central densities $\rho_{0}$, considering two representative values of the parameter, $\alpha=10^{-4}\alpha_{\star}$ and $\alpha=10^{-2}\alpha_{\star}$, as shown in Fig. 2.
@@
This allows us to provide a more comprehensive assessment of the effects of the non-minimal coupling term and the potential on stellar configurations with different central densities. It is evident that even when the parameter $\ell$ is as small as $\mathcal{O}(10^{-10})$, the mass and radius of neutron stars can still undergo significant modifications. Compared with the GR case, neutron stars exhibit smaller masses and radii at low central densities, while larger masses and radii are obtained in the high-central-density regime. Moreover, this effect becomes more pronounced for smaller values of the potential mass parameter $\alpha$. This behavior is similar to that observed in non-minimally coupled massive scalar-tensor gravity [29, 30, 31, 32, 40].
@@
Compared with the GR case, neutron stars in massive vector-tensor gravity exhibit smaller moments of inertia at low masses and larger moments of inertia at high masses. The influence of the non-minimal coupling term is relatively weak for low-mass stars but becomes more pronounced for high-mass stars. Moreover, this effect is stronger for smaller values of the parameter $\alpha$, indicating that the non-minimal coupling has a significant impact in the strong-field region, and that the magnitude of the vector-field mass plays an important role in this regime. This behavior is also similar to that observed in non-minimally coupled massive scalar-tensor gravity [29, 30, 31, 32, 40].
@@ V Conclusion
In this work, we have investigated neutron stars within a class of massive vector-tensor theories of gravity, characterized by a non-vanishing VEV of the vector field. Owing to the fact that the vector field selects a preferred spacetime direction, Lorentz symmetry is spontaneously broken in this theory. Exact static and spherically symmetric black-hole solutions had previously been constructed in this theory, for which condition (1) can be satisfied globally throughout spacetime. These solutions further enabled stringent Solar System constraints on the Lorentz-violating parameters of the theory. Motivated by this success, similar assumptions were subsequently adopted in attempts to construct compact star solutions. In Sec. III, however, we demonstrated that imposing such assumptions a priori inevitably leads to inconsistencies. We further showed that these assumptions are not fundamental requirements of the theory: instead they are dynamically enforced in the weak-field regime through the asymptotic boundary conditions at spatial infinity. As a result, the theory remains fully compatible with existing black hole solutions and Solar System tests, while providing a unified, natural, and self-consistent framework for constructing neutron star solutions.
@@
Based on our calculations of the neutron star mass, radius, and moment of inertia, we arrive at two main conclusions. First, even when the Lorentz-symmetry-breaking parameter $\ell$ is tightly constrained to very small values by Solar System tests, the presence of the vector-field potential term can still induce appreciable deviations in both the mass-radius relation and the moment of inertia of neutron stars relative to their GR counterparts. Second, although the explicit functional form of the potential does not affect black hole solutions, it has a significant impact on the properties of neutron stars. These features closely parallel those found in other massive non-minimally coupled gravity theories, such as massive scalar-tensor gravity.
@@
Finally, we outline several directions for future studies within this class of gravity theories. On the one hand, it is natural to investigate whether non-minimal gravitational couplings can have a significant impact on stellar stability [41]. A systematic analysis of the radial and non-radial oscillation stabilities and quasi-normal modes of neutron stars within this framework would be highly valuable. On the other hand, since the non-minimal coupling term and the potential term can affect the macroscopic properties of neutron stars, it is natural to consider additional observables, such as the quadrupole moment and tidal Love numbers. These quantities allow one to make use of approximately EOS-insensitive universal relations, including the I-Love-Q relations [42, 43], thereby providing complementary constraints on the Lorentz-violating parameter as well as on the form and parameters of the vector-field potential.
@@ Acknowledgements.
We are grateful to Peixiang Ji, H. Lü and Lijing Shao for useful discussions. S.L. and H.Y. were supported in part by the National Natural Science Foundation of China (No. 12105098, No. 12481540179, No. 12075084, No. 11690034, No. 11947216, and No. 12005059) and the Natural Science Foundation of Hunan Province (No. 2022JJ40264), and the innovative research group of Hunan Province under Grant No. 2024JJ1006, and by the Excellent Young Scholars Program of the Hunan Provincial Department of Education under Grant No. 25B0092.

### v2

@@ I Introduction
and by imposing Eq. (1) throughout the entire vacuum region [19]. Subsequent studies have further shown that Einstein-bumblebee gravity admits exact black hole solutions involving a richer set of physical parameters, including electric charge [20], magnetic charge [21], NUT charge [22], and cosmological constant [23]. This makes the theory a convenient and concrete framework for constraining spontaneous Lorentz symmetry breaking through astrophysical observations, as well as for investigating gravitationally induced quantum entanglement [24].
@@
Technically, there is no intrinsic difficulty in deriving compact star solutions from the Einstein-bumblebee gravity Lagrangian. Previous attempts failed primarily because of an unnecessarily strong and somewhat ad hoc theoretical assumption that condition (1) be imposed throughout the entire vacuum region. Einstein-bumblebee gravity, as a massive non-minimally coupled vector-tensor theory, is a higher-derivative gravity theory, for which vacuum spherically symmetric solutions are not unique. In this context, black hole solutions arise only as special vacuum solutions for which condition (1) can be consistently satisfied. For compact stars, however, different equations of state (EOSs) and central densities generically lead to different vacuum boundary conditions at the stellar surface. As a result, the exterior vacuum spacetime of a compact star may deviate from the equal-mass black hole geometry, and condition (1) is therefore no longer satisfied. Nevertheless, such deviations are confined mainly to the strong-field region, since both black hole and compact star solutions necessarily share the same boundary conditions at spatial infinity. This implies that assumption (1) need not be imposed a priori: it is naturally violated in the strong-field region while being dynamically enforced in the weak-field regime through asymptotic boundary conditions at spatial infinity. This treatment is well established in massive non-minimally coupled scalar-tensor gravity [30, 31, 32, 33, 34]. Consequently, consistency with previously obtained black hole solutions and existing observational constraints is preserved, and earlier Solar System tests of spontaneous Lorentz symmetry breaking based on exact black hole spacetimes remain physically meaningful.
@@ III.2 Vacuum equations
where the subscripts “in” and “ext” denote the interior and exterior solutions, respectively. The first derivatives of these functions are likewise continuous at $r=R$, although not written explicitly here. Since the curvature invariants can be expressed in terms of the metric functions and their derivatives, this continuity guarantees that the spacetime curvature remains regular across the stellar surface. It is worth noting that neutron-star spacetimes differ qualitatively from black-hole geometries in this respect. In certain black-hole solutions of this theory, the vector field can exhibit a coordinate singularity at the event horizon [21]. By contrast, in the neutron-star configurations considered here, the vector field remains regular throughout the entire spacetime, both in the interior and in the exterior regions.
We then continue the integration outward through the vacuum region up to spatial infinity. The vacuum field equations are given by Eqs. (12)–(13) with $p=0$ and $\rho=0$. The zeroth-order vacuum equations can be organized into a coupled system nonlinear ODEs, which take the schematic form
@@ IV Equilibrium configurations and moments of inertia
To verify the consistency of the neutron star solutions obtained in this work, we first plot the metric functions ($f,\phi^{-2},\rho,\lambda^{\prime},w/\Omega$), as functions of the radial coordinate $r$ in Fig. 1, for representative values of the parameters $\ell$ and $\alpha$.
Figure 1: The plots illustrate the numerical solutions of the functions ($f,\phi^{-2},\rho,\lambda^{\prime},w/\Omega$) in massive vector-tensor gravity with the parameters $\alpha=10^{-4}\alpha_{\star}$ (solid lines) and $\alpha=10^{-2}\alpha_{\star}$ (dashed lines). The parameter $\ell$ is fixed to $\ell=10^{-10}$. The central density is set to $\rho_{0}=1.2\times 10^{15}\textup{g/cm}^{3}$, and the SLy EOS is adopted.
Motivated by the bounds $(10^{-12},10^{-9})$ on the parameter $\ell$ inferred from three classical Solar System tests [19], we choose $\ell=10^{-10}$ for our numerical analysis. We consider two values of the parameter $\alpha$ namely $\alpha=10^{-4}\alpha_{\star}$ and $\alpha=10^{-2}\alpha_{\star}$. The central density is fixed at $\rho_{0}=1.2\times 10^{15}\textup{g/cm}^{3}$. For both choices of $\alpha$, the resulting neutron star mass is $1.51M_{\odot}$, while the corresponding radii are $7.48r_{\star}$ (11.0 km) and $7.58r_{\star}$ (11.1 km), and the moments of inertia are $34.92I_{\star}$ and $34.25I_{\star}$. From the upper panel of Fig. 1, one can clearly see that $\phi^{2}f$ deviates from unity both inside the neutron star and in the exterior strong-field region, while it approaches unity only in the weak-field regime and asymptotically at spatial infinity, as required by the boundary condition (24). Moreover, the deviation becomes more pronounced as the potential parameter $\alpha$ decreases.
Next, we plot the mass-radius MRM-R and mass-central density $M-\rho_{0}$ relations for the SLy EOS over a range of central densities $\rho_{0}$, considering two representative values of the parameter, $\alpha=10^{-4}\alpha_{\star}$ and $\alpha=10^{-2}\alpha_{\star}$, as shown in Fig. 2.
@@
This allows us to provide a more comprehensive assessment of the effects of the non-minimal coupling term and the potential on stellar configurations with different central densities. It is evident that even when the parameter $\ell$ is as small as $\mathcal{O}(10^{-10})$, the mass and radius of neutron stars can still undergo significant modifications. Compared with the GR case, neutron stars exhibit smaller masses and radii at low central densities, while larger masses and radii are obtained in the high-central-density regime. Overall, larger values of the vector-field mass parameter $\alpha$ drive the stellar macroscopic properties closer to the GR limit, whereas smaller values lead to more significant deviations from GR, approaching the behavior of the corresponding massless theory. This deviation from GR becomes increasingly pronounced as $\alpha$ decreases. Interestingly, the impact of the deviation depends on the central density: compared with GR, low-central-density neutron stars tend to exhibit smaller masses and radii, while high-central-density configurations can attain larger masses and radii. This qualitative trend is similar to that found in non-minimally coupled massive scalar-tensor gravity [30, 31, 32, 33, 34, 42].
@@
Compared with the GR case, neutron stars in massive vector-tensor gravity exhibit smaller moments of inertia at low masses and larger moments of inertia at high masses. The influence of the non-minimal coupling term is relatively weak for low-mass stars but becomes more pronounced for high-mass stars. Moreover, this effect is stronger for smaller values of the parameter $\alpha$, indicating that the non-minimal coupling has a significant impact in the strong-field region, and that the magnitude of the vector-field mass plays an important role in this regime. This behavior is also similar to that observed in non-minimally coupled massive scalar-tensor gravity [30, 31, 32, 33, 34, 42].
@@ V Conclusion
In this work, we have investigated neutron stars within a class of massive vector-tensor theories of gravity, characterized by a non-vanishing VEV of the vector field. Owing to the fact that the vector field selects a preferred spacetime direction, Lorentz symmetry is spontaneously broken in this theory. Exact static and spherically symmetric black-hole solutions had previously been constructed in this theory, for which condition (1) can be satisfied globally throughout spacetime. These solutions further enabled stringent Solar System constraints on the Lorentz-violating parameters of the theory. Motivated by this success, similar assumptions were subsequently adopted in attempts to construct compact star solutions. In Sec. III, however, we demonstrated that imposing such assumptions a priori inevitably leads to inconsistencies. We further showed that these assumptions are not fundamental requirements of the theory: instead they are dynamically enforced in the weak-field regime through the asymptotic boundary conditions at spatial infinity. As a result, the theory remains fully compatible with existing black hole solutions and Solar System tests, while providing a unified, natural, and self-consistent framework for constructing neutron star solutions. Notably, this theoretical observation is not limited to the construction of neutron star solutions in non-minimally coupled vector-tensor models of Lorentz-violating gravity, but can be extended to theories in which the metric is non-minimally coupled to additional antisymmetric two-form fields [25].
@@
Based on our calculations of the neutron star mass, radius, and moment of inertia, we arrive at two main conclusions. First, even when the Lorentz-symmetry-breaking parameter $\ell$ is tightly constrained to very small values by Solar System tests, the presence of the vector-field potential term can still induce appreciable deviations in both the mass-radius relation and the moment of inertia of neutron stars relative to their GR counterparts. Second, although the explicit functional form of the potential does not affect black hole solutions, it has a significant impact on the properties of neutron stars. More specifically, larger values of the vector-field mass parameter $\alpha$ drive the stellar macroscopic properties closer to the GR limit, whereas smaller values lead to greater deviations from GR, approaching the behavior of the corresponding massless theory. These features closely parallel those observed in massive scalar-tensor gravity in the Einstein frame, both with and without self-interaction potentials [30, 31, 32, 33, 34, 42]. A detailed comparison with neutron star solutions in Einstein-Kalb-Ramond gravity and in massive Brans-Dicke theory formulated in the Jordan frame with self-interaction potentials would therefore be worthwhile.
Finally, it is worth emphasizing that, although we have established the existence of neutron star solutions within a class of massive vector-tensor theories satisfying the asymptotic condition (24), such models—owing to the presence of self-interacting vector fields—may still exhibit dynamical pathologies [43, 44, 45, 46, 47, 48, 49, 50]. In particular, they can be prone to instabilities associated with loss of hyperbolicity and tachyonic modes. Several approaches have been proposed to mitigate the problem of dynamical evolution [50], and a more complete ultraviolet completion of the theory is expected to ultimately cure or regulate these difficulties. Nevertheless, a systematic understanding of how such dynamical instabilities can be consistently resolved remains a central open challenge in this class of gravity theories.
@@
Additional important avenues for future investigation include the following. First, it is natural to investigate whether non-minimal gravitational couplings can have a significant impact on stellar stability [51]. A systematic analysis of the radial and non-radial oscillation stabilities and quasi-normal modes of neutron stars within this framework would be highly valuable. Second, since the non-minimal coupling term and the potential term can affect the macroscopic properties of neutron stars, it is natural to consider additional observables, such as the quadrupole moment and tidal Love numbers. These quantities allow one to make use of approximately EOS-insensitive universal relations, including the I-Love-Q relations [52, 53], thereby providing complementary constraints on the Lorentz-violating parameter as well as on the form and parameters of the vector-field potential. Third, a promising direction is to investigate the impact of couplings between the self-interacting vector field and the Maxwell field on neutron-star structure. Recent studies indicate that nonlinear electrodynamics can nontrivially influence magnetar shape [54], suggesting that such interactions may likewise lead to novel phenomenological signatures.
@@ Acknowledgements.
We are grateful to the referee for the constructive and insightful comments on our paper. We are grateful to Peixiang Ji, H. Lü and Lijing Shao for useful discussions. S.L. and H.Y. were supported in part by the National Natural Science Foundation of China (No. 12105098, No. 12481540179, No. 12075084, No. 11690034, No. 11947216, and No. 12005059) and the Natural Science Foundation of Hunan Province (No. 2022JJ40264), and the innovative research group of Hunan Province under Grant No. 2024JJ1006, and by the Excellent Young Scholars Program of the Hunan Provincial Department of Education under Grant No. 25B0092.